# Towards Understanding Fine-Tuning Mechanisms of LLMs via Circuit Analysis

### v1

@@ 1 Introduction
Mechanistic Interpretability (MI) has become a powerful approach for exploring the inner workings of machine learning models, particularly Large Language Models (LLMs) Rai et al. (2024). It provides valuable insights into how information flows and transforms across different layers Ferrando et al. (2024). One of the most critical aspects of deploying LLMs in real-world scenarios is fine-tuning Chung et al. (2024). However, the interpretability of how pre-trained models improve during fine-tuning remains limited, and the underlying mechanisms enabling their success across tasks require further investigation.
@@
Many studies in MI regard models as computational graphs Geiger et al. (2021), where circuits are specific subgraphs that perform identifiable functions Wang et al. (2022). Notably, this framework has been successfully applied to various LLMs, revealing emergent behaviors within attention heads and Multi-Layer Perceptrons (MLPs) Heimersheim & Janiak (2023); Burns et al. (2023); Hanna et al. (2023); Gould et al. (2023). Moreover, circuits have recently been leveraged to investigate the fine-tuning process of language models, seeking to understand the mechanisms behind fine-tuning Prakash et al. (2024); Chhabra et al. (2024); Jain et al. (2024). However, these studies often focus on tasks where pre-trained models already perform well (e.g., GPT-2 Radford et al. (2019) achieves around 98% accuracy on the IOI task), or they use general data for fine-tuning rather than domain-specific datasets Prakash et al. (2024). Under such conditions, fine-tuning mainly enhances existing mechanisms (e.g., by adding some attention heads). Consequently, their arguments may not be applicable in more practical fine-tuning scenarios where models initially perform poorly and require fine-tuning on domain data.
@@
To better understand fine-tuning mechanisms in practical settings, it is crucial to focus on tasks where fine-tuning leads to performance improvements. In this work, we design a class of mathematical tasks on which pre-trained large language models initially perform poorly with low accuracy, yet demonstrates a performance boost after fine-tuned. We employ the Edge Attribution Patching with Integrated Gradients (EAP-IG) Hanna et al. (2024) method to identify circuits within both pre-trained and fine-tuned models. Surprisingly, we observe that this approach consistently finds circuits with high faithfulness, even though the two models differ markedly in performance (see §3). To further validate the stability of the discovered circuits, we introduce another circuit metric, robustness, which measures the stability of identified circuits by assessing their edge similarity under different perturbation ratios of the dataset. We show that when compared with a randomly initialized transformer model, the pre-trained model, despite exhibiting very low prediction accuracy, can still achieve substantially higher robustness. This finding further supports the validity of the circuits discovered during the fine-tuning process, irrespective of their prediction performance.
@@
Our Main Findings. Based on the circuits analysis techniques and tasks introduced in §3, we provide a comprehensive interpretation of the key factors in the fine-tuning process. Specifically, we focus on three central research questions and summarize our main observations as follows:
@@
    4) How do circuits evolve during the fine-tuning process? We use pythia-1.4B-deduped Biderman et al. (2023), gpt-neo-2.7B Black et al. (2021), opt-6.7B Zhang et al. (2022) to fine-tune on five math tasks. By extracting the circuits at each stage of the model during fine-tuning and analyzing these circuits, the circuits identified by EAP-IG demonstrate high fidelity in both pre-trained and fine-tuned models, despite significant performance differences. We observe that during fine-tuning, circuits gradually converge as modifications to nodes and edges decrease. Meanwhile, new circuits emerge after fine-tuning, with edge changes playing a more significant role in this process.
@@ 2.1 Mechanistic Interpretability
Mechanistic Interpretability investigates how components in large language models process and represent information Wang et al. (2024). At present, many MI studies have been applied in various fields of AI Safety. For instance, oversimplified probes risk Friedman et al. (2024), unlearning fabricated knowledge Sun et al. (2024), reducing toxicity via alignment Lee et al. (2024), mitigating hallucinations by editing representations Zhang et al. (2024), and generating truthful outputs through inference-time interventions Li et al. (2023). Other studies explore how local model edits propagate across tasks Cohen et al. (2024); Meng et al. (2023), Multi-Head Attention in-context learning Chen et al. (2024); Chen & Zou (2024) and enhance influence-function sampling Koh et al. . Specifically, our study examines how circuits evolve during fine-tuning for mathematical tasks, focusing on node and edge changes to reveal mechanisms behind performance improvements.
@@ 2.2 Circuit Analysis and Fine-Tuning
One direction of Circuit Analysis focuses on building complete circuits. Early work localizes factual associations in mid-layer modules Meng et al. (2022) and uses causal mediation to uncover biases Vig et al. (2020); Hase et al. (2023). Automated methods like Automated Circuit Discovery identify significant units Conmy et al. (2023), while techniques like attribution patching, and refine circuit extraction by handling near-zero gradients Syed et al. (2023); Hanna et al. (2024). Edge pruning Bhaskar et al. (2024) provide insights into building the edge of the circuit. Another line of research investigates the functional roles of circuit components, such as Attention heads Wu et al. (2024); McDougall et al. (2023); Olsson et al. (2022); Gould et al. (2023); Cabannes et al. (2024) and Feed Forward Networks (FFNs) / MLPs Geva et al. (2021; 2022); Bhattacharya & Bojar (2024). Additionally, circuits have been used to analyze specific tasks, such as factual knowledge retrieval Geva et al. (2023), arithmetic computation Stolfo et al. (2023), Greater Than task Hanna et al. (2023), and circuit recognition in Indirect Object Identification Wang et al. (2022). Unlike these analyses, which focus on smaller-scale tasks and models, our work offers a new lens on how circuits evolve specifically during fine-tuning on mathematical tasks, revealing crucial roles of edge changes.
As pre-trained language models scale, fine-tuning methods have emerged, optimizing only a small subset of parameters Ding et al. (2023). Parameter-efficient fine-tuning (PEFT) methods, such as LoRA Hu et al. (2021), reduce computational costs while preserving functionality Ding et al. (2023). Advances in LoRA, including pruning Zhou et al. (2024) and adaptive budget allocation Zhang et al. (2023); Liu et al. (2022); Lialin et al. (2024), further improve efficiency. In our study, we introduce a circuit-aware LoRA approach that adaptively assigns higher ranks to layers with more edge changes, boosting efficiency and accuracy in mathematical tasks, and further illustrates how combining circuits from subtasks can enhance performance in compositional tasks during fine-tuning.
@@ 3.1 Circuit Discovery: EAP-IG
Attribution patching is a technique for identifying circuits using two forward passes and one backward pass Syed et al. (2023). In our experiments, we use Edge Attribution Patching with Integrated Gradients (EAP-IG) Hanna et al. (2024), which addresses computational inefficiency in large models and resolves zero-gradient issues with KL divergence. EAP-IG computes importance scores by integrating gradients along the path between clean and corrupted activations, making it our method of choice. The formula for scoring each edge is:
@@
In this study, we choose $m=5$ based on Hanna et al.’s (2024) recommendations Hanna et al. (2024).
@@ 3.2 Circuit Evalutaion: Faithfulness and Robustness
Faithfulness. Faithfulness serves as a key metric to evaluate the reliability of circuits discovered in MI and it quantifies how closely a circuit replicates the behavior of the original model Wang et al. (2022); Chhabra et al. (2024); Prakash et al. (2024). We adopt Kullback-Leibler divergence (KL-divergence) as the metric, following Conmy et al. Conmy et al. (2023). Let $M$ denote the model and $C$ the discovered circuit. Faithfulness is defined as the percentage of the model’s performance captured by the circuit. The formula for faithfulness is:
@@
The robustness score is computed using the Jaccard Similarity Jaccard (1912) between the initial circuit $G_{0}$ and perturbed circuits $G_{p}$. The formula is:
@@ 3.3 Tasks Design
Function Evaluation (Function). This task focuses on the model’s ability to compute values for linear functions, typically of the form $y=mx+b$. Corrupted data involves altering the constant term in the function. The dataset contains 5,000 instances. For each task, we ensure a strict separation between the dataset used for fine-tuning and the dataset used for circuit analysis. Specifically, 80% of the dataset is allocated for fine-tuning, and the remaining 20% is reserved for identifying circuits and evaluating the model’s and circuit’s accuracies. This separation guarantees that performance evaluation is conducted on data unseen during fine-tuning.
@@ 4.1 Model Accuracy, Circuit Faithfulness, and Robustness Analysis
To analyze circuit evolution, we first evaluate model accuracy across fine-tuning checkpoints. We use LoRA Hu et al. (2021) to fine-tune the Pythia-1.4B model  Biderman et al. (2023) on five different mathematical tasks. The experimental settings for fine-tuning are shown in Appendix A. The left panel of Figure 2 depicts the accuracy dynamics of the model on five mathematical tasks during fine-tuning. We track the model’s accuracy at various training stages across different tasks, revealing consistent improvements in performance throughout the fine-tuning process.
Next, we explore the faithfulness of the circuits found at each stage of fine-tuning. Prior work Hanna et al. (2024) achieved over 85% faithfulness by selecting 1–2% of edges. Given our more complex tasks and larger model, we select 5% of edges to ensure reliable circuits (faithfulness >70%). As shown in the middle panel of Figure 2, circuit faithfulness consistently exceeds 80% across most tasks, both before fine-tuning (Checkpoint 0) and throughout fine-tuning (Checkpoints 1–10). The only exception is Add/Sub task, where faithfulness is 77.52% before fine-tuning. These results confirm high circuit faithfulness in both pre-trained and fine-tuned models across all tasks.
@@ 4.3 Reorganizing Circuit Edges to Form a New Circuit
As discussed in Section 3.1, each edge’s score is computed as the dot product of the averaged loss gradients and activation difference, quantifying its influence on model predictions. To examine structural changes in circuits during fine-tuning, we use the 95th percentile of edge scores as a dynamic threshold. Edges in the initial and final circuits exceeding this threshold are retained, yielding sparser circuits that capture the model’s core information flow. Experimental results for all other tasks are provided in Appendix E. The distribution of added and deleted nodes and edges follows a distinct pattern. As illustrated in the left part of Figure 3, added nodes are predominantly located in the middle and later layers of the circuit, whereas added and deleted edges are concentrated in the middle layers. The shallow layers exhibit minimal changes, providing a stable foundation for task-specific adaptations.
@@ 5.2 Improving Fine-Tuning Efficiency and Accuracy by Circuit Insights
Table 1: Performance metrics for Add/Sub (within 300) and four other math tasks: Mul/Div, Sequence, LCM, and Function across different configurations. The control groups of CircuitLoRA ($r_{o}=8$, $r_{c}=32$) are LoRA ($r_{o}=16$) and RandomLoRA ($r_{o}=8$, $r_{c}=32$), and the control groups of CircuitLoRA ($r_{o}=32$, $r_{c}=64$) are LoRA ($r_{o}=32$) and RandomLoRA ($r_{o}=32$, $r_{c}=64$). Here, $r_{o}$ and $r_{c}$ represent the ranks used in CircuitLoRA, where $r_{c}$ is the rank for critical layer modules, and $r_{o}$ is the rank for non-critical layer modules. Model Accuracy is expressed as percentages.
@@ 6.1 Compositional Tasks, Compositional Circuits and Union Circuits
Compositional Tasks. A compositional task is composed of multiple interrelated subtasks, each contributing to the overall objective. Compositional Circuits. A Compositional Circuit captures the structural and functional dependencies within a compositional task. Union Circuits. A Union Circuit is constructed by merging the circuits of all subtasks.
To design the compositional tasks, we consider the two-step operation, which involves the calculation of two different types of mathematical problem, such as addition/subtraction and multiplication/division. For instance, the compositional task ‘$(61-45)\times 45=$” involves two mathematical operations: (1) (Addition/Subtraction): $61-45=$”; and (2) (Multiplication/Division): $16\times 45=$”. More examples of compositional tasks can be found in Appendix J.
@@ 6.2 Efficient Single-Phase Fine-Tuning on Compositional Task with Union Circuit
Then, we use Union Circuit and Compositional Circuit to identify the Critical Layers to further explore the ‘approximation ability” of Union Circuit. Table 3 summarizes the performance of CircuitLoRA and LoRA on the two-step operations task. Specifically, CircuitLoRA with Compositional Circuit achieves the highest accuracy of 67.20%. Surprisingly, when using the Union Circuit for Critical Layer identification, CircuitLoRA achieves 65.50%, still exceeding the performance of LoRA except the Compositional Circuit configuration.

### v2

@@ 1 Introduction
Mechanistic Interpretability (MI) has become a powerful approach for exploring the inner workings of machine learning models, particularly Large Language Models (LLMs) (Rai et al., 2024). It provides valuable insights into how information flows and transforms across different layers (Ferrando et al., 2024). One of the most critical aspects of deploying LLMs in real-world scenarios is fine-tuning (Chung et al., 2024). However, the interpretability of how pre-trained models improve during fine-tuning remains limited, and the underlying mechanisms enabling their success across tasks require further investigation.
@@
Many studies in MI regard models as computational graphs (Geiger et al., 2021), where circuits are specific subgraphs that perform identifiable functions (Wang et al., 2022). Notably, this framework has been successfully applied to various LLMs, revealing emergent behaviors within attention heads and Multi-Layer Perceptrons (MLPs) (Heimersheim & Janiak, 2023; Burns et al., 2023; Hanna et al., 2023; Gould et al., 2023). Moreover, circuits have recently been leveraged to investigate the fine-tuning process of language models, seeking to understand the mechanisms behind fine-tuning (Prakash et al., 2024; Chhabra et al., 2024; Jain et al., 2024). However, these studies often focus on tasks where pre-trained models already perform well (e.g., GPT-2 (Radford et al., 2019) achieves around 98% accuracy on the IOI task), or they use general data for fine-tuning rather than domain-specific datasets (Prakash et al., 2024). Under such conditions, fine-tuning mainly enhances existing mechanisms (e.g., by adding some attention heads). Consequently, their arguments may not be applicable in more practical fine-tuning scenarios where models initially perform poorly and require fine-tuning on domain data.
@@
To better understand fine-tuning mechanisms in practical settings, it is crucial to focus on tasks where fine-tuning leads to performance improvements. In this work, we design a class of mathematical tasks on which pre-trained large language models initially perform poorly with low accuracy, yet demonstrates a performance boost after fine-tuned. We employ the Edge Attribution Patching with Integrated Gradients (EAP-IG) (Hanna et al., 2024) method to identify circuits within both pre-trained and fine-tuned models. Surprisingly, we observe that this approach consistently finds circuits with high faithfulness, even though the two models differ markedly in performance (see §3). To further validate the stability of the discovered circuits, we introduce another circuit metric, robustness, which measures the stability of identified circuits by assessing their edge similarity under different perturbation ratios of the dataset. We show that when compared with a randomly initialized transformer model, the pre-trained model, despite exhibiting very low prediction accuracy, can still achieve substantially higher robustness. This finding further supports the validity of the circuits discovered during the fine-tuning process, irrespective of their prediction performance.
@@
Our Main Findings. Based on the circuits analysis techniques and tasks introduced in §3, we provide a comprehensive interpretation of the key factors in the fine-tuning process. Specifically, we focus on three central research questions and summarize our main observations as follows. The code and data are available at https://github.com/Xu0615/FinetuneCircuits.
@@
    4) How do circuits evolve during the fine-tuning process? We use pythia-1.4B-deduped (Biderman et al., 2023), gpt-neo-2.7B (Black et al., 2021), opt-6.7B (Zhang et al., 2022) to fine-tune on five math tasks. By extracting the circuits at each stage of the model during fine-tuning and analyzing these circuits, the circuits identified by EAP-IG demonstrate high fidelity in both pre-trained and fine-tuned models, despite significant performance differences. We observe that during fine-tuning, circuits gradually converge as modifications to nodes and edges decrease. Meanwhile, new circuits emerge after fine-tuning, with edge changes playing a more significant role in this process.
@@ 2.1 Mechanistic Interpretability
Mechanistic Interpretability investigates how components in large language models process and represent information (Wang et al., 2024). At present, many MI studies have been applied in various fields of AI Safety. For instance, oversimplified probes risk (Friedman et al., 2024), unlearning fabricated knowledge (Sun et al., 2024), reducing toxicity via alignment (Lee et al., 2024), mitigating hallucinations by editing representations (Zhang et al., 2024), and generating truthful outputs through inference-time interventions (Li et al., 2023). Other studies explore how local model edits propagate across tasks (Cohen et al., 2024; Meng et al., 2023), Multi-Head Attention in-context learning (Chen et al., 2024; Chen & Zou, 2024) and enhance influence-function sampling (Koh et al., ). Specifically, our study examines how circuits evolve during fine-tuning for mathematical tasks, focusing on node and edge changes to reveal mechanisms behind performance improvements.
@@ 2.2 Circuit Analysis and Fine-Tuning
One direction of Circuit Analysis focuses on building complete circuits. Early work localizes factual associations in mid-layer modules (Meng et al., 2022) and uses causal mediation to uncover biases (Vig et al., 2020; Hase et al., 2023). Automated methods like Automated Circuit Discovery identify significant units (Conmy et al., 2023), while techniques like attribution patching, and refine circuit extraction by handling near-zero gradients (Syed et al., 2023; Hanna et al., 2024). Edge pruning (Bhaskar et al., 2024) provide insights into building the edge of the circuit. Another line of research investigates the functional roles of circuit components, such as Attention heads (Wu et al., 2024; McDougall et al., 2023; Olsson et al., 2022; Gould et al., 2023; Cabannes et al., 2024) and Feed Forward Networks (FFNs) / MLPs (Geva et al., 2021, 2022; Bhattacharya & Bojar, 2024). Additionally, circuits have been used to analyze specific tasks, such as factual knowledge retrieval (Geva et al., 2023), arithmetic computation (Stolfo et al., 2023), Greater Than task (Hanna et al., 2023), and circuit recognition in Indirect Object Identification (Wang et al., 2022). Unlike these analyses, which focus on smaller-scale tasks and models, our work offers a new lens on how circuits evolve specifically during fine-tuning on mathematical tasks, revealing crucial roles of edge changes.
As pre-trained language models scale, fine-tuning methods have emerged, optimizing only a small subset of parameters (Ding et al., 2023). Parameter-efficient fine-tuning (PEFT) methods, such as LoRA (Hu et al., 2021), reduce computational costs while preserving functionality (Ding et al., 2023). Advances in LoRA, including pruning (Zhou et al., 2024) and adaptive budget allocation (Zhang et al., 2023; Liu et al., 2022; Lialin et al., 2024), further improve efficiency. In our study, we introduce a circuit-aware LoRA approach that adaptively assigns higher ranks to layers with more edge changes, boosting efficiency and accuracy in mathematical tasks, and further illustrates how combining circuits from subtasks can enhance performance in compositional tasks during fine-tuning.
@@ 3.1 Circuit Discovery: EAP-IG
Attribution patching is a technique for identifying circuits using two forward passes and one backward pass (Syed et al., 2023). In our experiments, we use Edge Attribution Patching with Integrated Gradients (EAP-IG) (Hanna et al., 2024), which addresses computational inefficiency in large models and resolves zero-gradient issues with KL divergence. EAP-IG computes importance scores by integrating gradients along the path between clean and corrupted activations, making it our method of choice. The formula for scoring each edge is:
@@
In this study, we choose $m=5$ based on Hanna et al.’s (2024) recommendations (Hanna et al., 2024).
@@ 3.2 Circuit Evalutaion: Faithfulness and Robustness
Faithfulness. Faithfulness serves as a key metric to evaluate the reliability of circuits discovered in MI and it quantifies how closely a circuit replicates the behavior of the original model (Wang et al., 2022; Chhabra et al., 2024; Prakash et al., 2024). We adopt Kullback-Leibler divergence (KL-divergence) as the metric, following Conmy et al. (Conmy et al., 2023). Let $M$ denote the model and $C$ the discovered circuit. Faithfulness is defined as the percentage of the model’s performance captured by the circuit. The formula for faithfulness is:
@@
The robustness score is computed using the Jaccard Similarity (Jaccard, 1912) between the initial circuit $G_{0}$ and perturbed circuits $G_{p}$. The formula is:
@@ 3.3 Tasks Design
Function Evaluation (Function). This task focuses on the model’s ability to compute values for linear functions, typically of the form $y=mx+b$. Corrupted data involves altering the constant term in the function. The dataset contains 5,000 instances.
For each task, we ensure a strict separation between the dataset used for fine-tuning and the dataset used for circuit analysis. Specifically, 80% of the dataset is allocated for fine-tuning, and the remaining 20% is reserved for identifying circuits and evaluating the model’s and circuit’s accuracies. This separation guarantees that performance evaluation is conducted on data unseen during fine-tuning.
@@ 4.1 Model Accuracy, Circuit Faithfulness, and Robustness Analysis
To analyze circuit evolution, we first evaluate model accuracy across fine-tuning checkpoints. We use LoRA (Hu et al., 2021) to fine-tune the Pythia-1.4B model  (Biderman et al., 2023) on five different mathematical tasks. The experimental settings for fine-tuning are shown in Appendix A. The left panel of Figure 2 depicts the accuracy dynamics of the model on five mathematical tasks during fine-tuning. We track the model’s accuracy at various training stages across different tasks, revealing consistent improvements in performance throughout the fine-tuning process.
Next, we explore the faithfulness of the circuits found at each stage of fine-tuning. Prior work (Hanna et al., 2024) achieved over 85% faithfulness by selecting 1–2% of edges. Given our more complex tasks and larger model, we select 5% of edges to ensure reliable circuits (faithfulness >70%). As shown in the middle panel of Figure 2, circuit faithfulness consistently exceeds 80% across most tasks, both before fine-tuning (Checkpoint 0) and throughout fine-tuning (Checkpoints 1–10). The only exception is Add/Sub task, where faithfulness is 77.52% before fine-tuning. These results confirm high circuit faithfulness in both pre-trained and fine-tuned models across all tasks.
@@ 4.3 Reorganizing Circuit Edges to Form a New Circuit
As discussed in Section 3.1, each edge’s score is computed as the dot product of the averaged loss gradients and activation difference, quantifying its influence on model predictions. To examine structural changes in circuits during fine-tuning, we use the 95th percentile of edge scores as a dynamic threshold. Edges in the initial and final circuits exceeding this threshold are retained, yielding sparser circuits that capture the model’s core information flow. Experimental results for all other tasks are provided in Appendix E.
The distribution of added and deleted nodes and edges follows a distinct pattern. As illustrated in the left part of Figure 3, added nodes are predominantly located in the middle and later layers of the circuit, whereas added and deleted edges are concentrated in the middle layers. The shallow layers exhibit minimal changes, providing a stable foundation for task-specific adaptations.
@@ 5.2 Improving Fine-Tuning Efficiency and Accuracy by Circuit Insights
Table 1: Performance metrics for Add/Sub (within 300) and four other math tasks: Mul/Div, Sequence, LCM, and Function across different configurations. The control groups of CircuitLoRA ($r_{o}=8$, $r_{c}=32$) are LoRA ($r_{o}=16$) and RandomLoRA ($r_{o}=8$, $r_{c}=32$), and the control groups of CircuitLoRA ($r_{o}=16$, $r_{c}=64$) are LoRA ($r_{o}=32$) and RandomLoRA ($r_{o}=16$, $r_{c}=64$). Here, $r_{o}$ and $r_{c}$ represent the ranks used in CircuitLoRA, where $r_{c}$ is the rank for critical layer modules, and $r_{o}$ is the rank for non-critical layer modules. Model Accuracy is expressed as percentages.
@@ 6.1 Compositional Tasks, Compositional Circuits and Union Circuits
Compositional Tasks. A compositional task consists of a sequence or combination of two or more simpler subtasks, where the output of one subtask often serves as the input to the next. For example, computing $(61-45)\times 45$ first requires solving the subtraction $(61-45)$, then using its result in a multiplication. By breaking complex reasoning into these interrelated steps, we can isolate and analyze each module’s contribution to overall performance.
Compositional Circuits. A Compositional Circuit is the end-to-end subnetwork of the model that directly implements a compositional task. It captures both the intra-subtask pathways and the cross-step dependencies that arise when information must flow from one operation (e.g. subtraction) into the next (e.g. multiplication). Extracting this circuit requires running the discovery pipeline on the full compositional task.
Union Circuits. A Union Circuit is formed by taking the edge-union of individual subtask circuits without re-extracting a dedicated compositional circuit. By merging the critical edges and nodes from each primitive operation’s circuit—while preserving edge counts for fair comparison—the Union Circuit approximates the full Compositional Circuit at a fraction of the discovery cost.
To design the compositional tasks, we consider the two-step operation, which involves the calculation of two different types of mathematical problem, such as addition/subtraction and multiplication/division. For instance, the compositional task “$(61-45)\times 45=$” involves two mathematical operations: (1) (Addition/Subtraction): “$61-45=$”; and (2) (Multiplication/Division): “$16\times 45=$”. More examples of compositional tasks can be found in Appendix J.
@@ 6.2 Efficient Single-Phase Fine-Tuning on Compositional Task with Union Circuit
Then, we use Union Circuit and Compositional Circuit to identify the Critical Layers to further explore the “approximation ability” of Union Circuit. Table 3 summarizes the performance of CircuitLoRA and LoRA on the two-step operations task. Specifically, CircuitLoRA with Compositional Circuit achieves the highest accuracy of 67.20%. Surprisingly, when using the Union Circuit for Critical Layer identification, CircuitLoRA achieves 65.50%, still exceeding the performance of LoRA except the Compositional Circuit configuration.